Tags:
create new tag
view all tags

20101111 Shift

09:00 - 25:00 AA, LC, LD, LN

Plan

Take access to

  • Align laser
  • Check Apd Dump - apt stages not registering
  • Change over to laserwire from OTR setup
  • Try OTR camera again
  • Reset / test energy meter with laser

  • Load Optics
  • Tune for low background
  • Wire scan for angle
  • Adjust Angle
  • Iterate for angle / background
  • Laser timing
  • Find location of e-beam with OTR screen
  • Try for collisions

Start (09:00)

LC found seed laser not locked this morning (having been on over night). Wouldn't lock. Adjusted manually thumb screw in laser to adjust frequency and looked at difference frequency on lock box. Laser power down to 300mW.

Moving access back until laser ready so we take fewer accesses.

AA found beam was gone entirely at QD4AFF. Used ZV1FF ZH1FF ZH10X ZV11X (between extraction line and final focus section) to get beam back. Background back to normal.

AA tuned end of linac to get more charge into transport line.

Tuning (10:00)

AA LN loaded set10nov10_0004.dat for extraction line quadrupoles (all).

AA LN tuned steering magnets pre LWIP.

Adjusted attenuation on cherenkov detector down from 31dB to 12dB. Signal around 1V / 10000ADC Counts.

When ADC saturates it stops reading out and at the same time the OTR screen & Chamber stopped responding - must check.

LD tuning 11:00 - 13:00 - notes here

Finished with background around 7000 ADC counts (~700mV with new attenuation). Check wire scanners - sure enough one was clipping the beam - background down to around 5000.

Wire Scan - Orbit (13:00)

Started wire scan, but wire scanner won't move correctly. AA / Company guys fixed it.

LC check all wire scanners - as usual, they were clipping the beam. By checking and moving all the wires clear of the beam. Background dropped to ~1200 ADC Counts - which was 125mV!!! Really! (Before the attenuator).

Wire Scanner Dimension Centre Width Data File
MW1X X 20.515 159.626.3 lwLog_20101111_135825.dat
Y 55.084 14.91.7 lwLog_20101111_140737.dat
MW2X X 34.191 129.514.5 lwLog_20101111_142118.dat
Y 64.692 29.57.3 lwLog20101111_143327.dat

Set file - set10nov11_1414.dat

Orbit pointing upwards a little bit but if laser was centred, should be just within window. Horizontally it was spot on.

Noticed BPM array in ExtLwProcess.py was wrong way round. Should be 21, 20, 19, 18. Fixed. Our EDM display now actually matches the BPM display.

Access 14:00 - 19:00

Aligned Laser through transport line, through IP with chamber at previous position and checked back reflection along transport line.

Switched from OTR post IP to Laserwire PostIP. AA tried OTR camera - still dead.

Tried energy meter with laser on it self triggered - didn't work. Doesn't work triggered either. Tried handheld meter with screen, but had wrong firmware and didn't recognise drivers. Went back to SLink original energy meter and use the Gentec software to read out on monitor. Won't give energy in EPICS but will do on monitor so we know if there's anything wrong by a sudden drop in energy.

LC cleaned FF lens which was quite dirty. Was in dry fridge over summer, but perhaps something else was in there too and the humidity left a residue. Cleaned though.

Checked energy loss along transport line - not bad at all (~20%).

Checked APD dump stage.

Checked APD Signal with laser.

Resume

AA started beam again at 19:00. Tuning for transmission through whole accelerator. Background around 1500ADC Counts. Charge around 0.4 10^10.

Final wire scan before searching for collisions.

Wire Scanner Dimension Centre Width
MW1X X 20.493 169.631.2
Y 54.126 13.62.0
MW2X X 34.259 115.417.1
Y 63.011 27.25.5

Delta Theta x = 28 μrad Delta Theta y = 6 μrad

But see note on corrector magnet ZV11X

Edit | Attach | Watch | Print version | History: r8 < r7 < r6 < r5 < r4 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r8 - 21 Nov 2010 - LauraCorner

 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback